OpenStudy (anonymous):

Combine into a single logarithm. 3log (x+y) + 2log(x-y) - log(x^2+y^2)

OpenStudy (kropot72):

As a first step the coefficients of the first two terms can be eliminated giving \[\log_{} (x+y)^{3}+\log_{} (x-y)^{2}-\log_{} (x ^{2}+y ^{2})\] Do you understand this step?

OpenStudy (anonymous):

I understand.

OpenStudy (kropot72):

The next step is to use the properties of logs \[\log a + \log b - \log c=\log\frac{a \times b}{c}\]